### v3

@@ 2.4.3 The Maximum Likelihood Decision Rule Maximizes Entropy
We can also motivate our maximum likelihood rule through the principle of maximum entropy (Jaynes, 1957a, b), which unites the theory of information with statistical physics. In Section 1.1, we discuss how differences in entropy drive variation in the likelihood such that the maximum likelihood distribution is also the maximum entropy distribution. Jaynes (1968) advocates for application of the principle of maximum entropy to circumvent the subjective specification of a prior. To make Bayesian decision-making more objective, one option is to choose the least informative prior. Jaynes’ alternative to make the process more objective is to choose the least informative updated distribution. The distribution that can generate the data in the greatest number of ways—the distribution that maximizes entropy, also the distribution that maximizes our design-based likelihood—is the least informative updated distribution. In Appendix F, we illustrate with a figure similar to Figure 1 that in the Monty Hall game, a celebrated application of Bayesian decision making, it is possible to make the same decision without specifying a Bayesian prior using a maximum likelihood decision rule and the principle of maximum entropy.
3 Applications: Counting Defiers
@@
We apply our rule to two published experiments with positive, statistically significant average effects on takeup of desired health behaviors and plausible defiers. Figure 3 summarizes the results from all possible empirical applications in experiments with given sizes, illustrating empirical conditions under which our rule yields defiers. It is predictable from the data for each application and the patterns in Figure 3 that our first application yields no defiers, and our second application yields defiers, which we confirm with exhaustive grid search. Our focus here is to provide examples for applied researchers to follow when presenting and interpreting quantitative results and to provide substantive insights. The top of Table 1 summarizes the context and design for both applications.
@@
In Tappin et al. (2015), the sample includes 612 pregnant smokers, and the definition of “takeup” is to quit smoking, which is a desired health behavior given concerns that smoking can cause pregnancy complications and lasting adverse effects on children. The experiment uses a completely randomized design, with half of the subjects assigned to an intervention arm with a payment and the other half assigned to a control arm with usual care. Monotonicity assumptions have a long history in the context of interventions that induce pregnant women to quit smoking. In a study discussed by Angrist, Imbens, and Rubin (1996), Permutt and Hebel (1989) proposed an early monotonicity assumption requiring that women who would quit smoking in control would also quit smoking in an arm that received a multi-faceted intervention. However, literature in psychology and economics expresses concern that payments can backfire in many contexts, including on-time pickup of children from day care (Gneezy and Rustichini, 2000), blood donation (Mellström and Johannesson, 2008; Lacetera and Macis, 2010), and vaccination (Schneider, Campos-Mercade, Meier, Pope, Wengström, and Meier, 2023). Defiers are plausible in Tappin et al. (2015) because a payment could backfire by crowding out pregnant women’s intrinsic motivations to quit smoking based on moral ideals and health goals.
Table 1: Applications: Standard Statistics and Proposed Design-Based Maximum Likelihood Estimates
@@
The average effect of the payment relative to usual care is a statistically significant 14 percentage point increase in the smoking quit rate on a base of 8% in control. Our proposed decision rule yields evidence beyond the average effect: our maximum likelihood estimate includes $52/612=8\%$ always takers (effect=0), $86/612=14\%$ compliers (effect=1), and $474/612=77\%$ never takers (effect=0). The weighted average effect across the three types ($8\times 0+14\times 1+77\times 0$) is equal to the estimated average effect of 14 percentage points, which is also equal to the share of compliers. We would obtain the same results if we assumed monotonicity in the sample, so our rule provides evidence in favor of monotonicity without assuming it.
@@
We emphasize that even though our proposed decision rule provides evidence that there are no defiers in the sample, there is considerable uncertainty in our maximum likelihood estimate. As discussed in Appendix E, we report a 95% smallest credible set on the number of defiers in the sample, which we construct by reporting the collection of numbers of defiers occurring within the 95% smallest credible set on the joint distribution of potential outcomes in the sample. We caution that because the numbers of always takers, compliers, defiers, and never takers must sum to the sample size, the credible sets for each of the four types are dependent. The smallest credible set on defiers includes up to 71 defiers, which represents 71/612=12% of the sample size.
@@
Building on the evidence provided by our decision rule, researchers could go on to make the stronger assumption of monotonicity and proceed as usual. With additional data on covariates, they could calculate average characteristics of always takers, compliers, and never takers (Imbens and Rubin, 1997; Katz, Kling, Liebman, et al., 2001; Abadie, 2002, 2003). With additional data on a second stage outcome like birth weight, they could interpret the resulting instrumental variable estimate as a local average treatment effect on compliers (Imbens and Angrist, 1994).
@@
Our second example, Johnson and Goldstein (2003), is a high profile experiment in which a simple change from an opt-in default to an opt-out default increases organ donation almost twofold from 43% in control to 82% in intervention. Results from Johnson and Goldstein (2003) have inspired state policy changes to encourage life-saving organ donations to combat chronic organ shortages (Kessler and Roth, 2025). However, in a recent experiment by Kessler and Roth (2025), in which takeup represents actually becoming an organ donor instead of hypothetically becoming an organ donor, the average effect is close to zero. Especially given the hypothetical outcome in Johnson and Goldstein (2003), both compliers and defiers are plausible. Compliers may think of the default as an endorsement by the experimenter and thus follow it regardless of what it is. Defiers may think of a default as a threat to their autonomy and thus go against it regardless of what it is.
@@
Given the plausibility of defiers in this context, a monotonicity assumption might not be reasonable, but there is no obvious alternative, so standard practice is to engage with bounds on defiers, which can be large. As shown at the bottom of Table A.1, in which we report optional auxiliary statistics, the smallest possible number of defiers in the sample is 0 because it is always possible for all people who take up to be always takers and all people who do not take up to be never takers, as in the null of Fisher’s exact test. The largest possible number of defiers in the sample occurs if all 11(=6150)11(=61-50) people who do not take up in intervention and all 23 people who take up in control are defiers, yielding a maximum defier share of $30\%(=(23+11)/115$). The estimated Fréchet bounds are tighter because they impose that the estimated marginal distribution of potential outcomes in one arm yields the true marginal distribution of potential outcomes in the other arm and thus in the sample. To calculate the upper Fréchet bound from (3), we consider the bounds on the share of defiers implied by the takeup rates in intervention and control. The intervention takeup rate implies that the share of defiers in intervention and thus in the sample is at most $18\%(=100\%-82\%)$, and the control takeup rate implies that the share of defiers in control and thus in the sample is at most 43%, so the share in intervention binds, and the upper Fréchet bound is 18%.
@@
The Fréchet bounds are estimated with error, so there is even greater uncertainty in the number of defiers. The maximum number of defiers in our 95% smallest credible set is 34, which represents 30% of the sample of 115. The 95% confidence interval on the population fraction of defiers from Imbens and Angrist (1994) also accounts for estimation of the Fréchet bounds, so we report it for comparison. It is slightly narrower, extending from 0% to 26%, but still includes the estimated Fréchet bounds. Accounting for estimation of the Fréchet bounds is empirically important. As demonstrated in Figure 2, if we do not account for estimation of the Fréchet bounds and instead construct a 95% smallest credible set on defiers within the estimated Fréchet set, variation in the likelihood allows us to exclude 8 and 9 defiers—counts in the middle of the Fréchet bounds translated into whole numbers from 0 to 21.
@@
In this application, evidence from the design-based model suggests a weaker alternative assumption to monotonicity—an assumption of no never takers—that can subsequently be imposed. While the assumption of no never takers is still strong, by the principle of maximum entropy, it “assumes the least” about the joint distribution of potential outcomes (Jaynes, 1968). From the design-based likelihood in (2), the assumption of no never takers can generate the data in $(28\text{ choose }15)\times(66\text{ choose }35)\times(21\text{ choose }11)=8.5\times 10^{31}$ randomized assignments representing 3.4% of the total randomized assignments ($(115\text{ choose }61)=2.5\times 10^{33}$). The assumption of no defiers can generate the data in only $(49\text{ choose }26)\times(45\text{ choose }24)\times(21\text{ choose }11)=7.8\times 10^{31}$ randomized assignments, representing 3.1% of the total randomized assignments. An ex post justification for the assumption of no never takers is that though experimental subjects might feel comfortable complying with or defying defaults imposed by the experimenter, they might feel uncomfortable revealing to the experimenter that they would not donate their organs under either default, especially since they do not know when they are asked a hypothetical question if a later question will ask them for a response under the other default.
@@
If a second stage outcome such as actual organ donation is available, under an assumption of no never takers, the instrumental variable estimate is no longer interpretable as an average treatment effect on compliers without further assumptions. As discussed by Angrist, Imbens, and Rubin (1996), under an additional assumption that average treatment effects are the same for compliers and defiers, the instrumental variable estimate is again interpretable as a LATE. Tchetgen Tchetgen (2024) proposes an alternative assumption to interpret the instrumental variable estimand in the presence of compliers and defiers as a “nudge average treatment effect.”
@@
The value of an assumption of no never takers becomes apparent if data are available on covariates. Under the assumption of no never takers, it is possible to label some specific people as compilers and others as defiers: all people in intervention who do not take up must be compliers, and all people who do not take up in control must be defiers. This reasoning allows us to make claims that the intervention had a causal effect on some specific people in the sample. Using terminology from Pearl (1999), under an assumption of no never takers, the intervention was a “necessary” cause of no takeup for all 11 people assigned intervention who do not take up, and was is a “sufficient” cause of no takeup for all 31 people assigned control who do not take up. Under a monotonicity assumption, it is not possible to directly label any specific people as compliers or defiers because people who take up in intervention can be compliers or always takers, and people who do not take up in control can be compliers and never takers. An assumption that allows us to label some specific people as compilers and others as defiers could be an important first step to learn their characteristics, improve targeting in future samples, and obtain a larger average effect.
@@ 4 Contributions to the Literature
Our primary insight is that a randomized experiment can yield evidence on an estimand more primitive than the average effect: the joint distribution of potential outcomes in the sample. The evidence can support a monotonicity assumption or inform a specific alternative assumption. We obtain evidence by using the structure of the implemented randomization process to construct a design-based model and its resulting likelihood function, and we evaluate this evidence in a statistical decision theory framework. Our work contributes to the literature on design-based econometrics, statistical decision theory, and causal inference without monotonicity.
@@
In his 1935 book on the design of experiments, Fisher notes that the design of Charles Darwin’s experiments was “greatly superior” to the “statistical methods available at the time” (Fisher, 1935). In their 2017 handbook chapter on the econometrics of experiments, Susan Athey and Guido Imbens write that they “recommend using statistical methods that are directly justified by randomization, in contrast to the more traditional sampling-based approach that is commonly used in econometrics” (Athey and Imbens, 2017). The sampling-based approach assumes that people in an experiment are sampled randomly from a population, and it attempts to learn about the population. The design-based approach, which we adopt, attempts to learn about the sample of people in the experiment, which is useful since participation in experiments and clinical trials can be far from random (Alsan, Cawley, Doyle, and Skelley, 2025) and some samples do not come from a meaningful population (Abadie, Athey, Imbens, and Wooldridge, 2020).
@@
The design-based model enables our advances to the frontier. Our ability to learn about defiers with a design-based likelihood may seem paradoxical because it is well known that the traditional sampling-based likelihood does not vary with the number of defiers within the Fréchet bounds determined by the estimated marginal distributions of potential outcomes in intervention and control. As we review in Appendix C, for a given pair of marginal distributions, the likelihood function in the traditional sampling-based model is flat over all copulas connecting them. A large literature has focused on specifying what we can learn from the Fréchet bounds in the traditional sampling-based model (Balke and Pearl 1997; Heckman, Smith, and Clements 1997; Manski 1997a; Tian and Pearl 2000; Zhang and Rubin 2003; Imbens and Manski 2004; Fan and Park 2010; Mullahy 2018; Ding and Miratrix 2019; Li and Pearl 2019; Bai, Huang, Moon, Shaikh, and Vytlacil 2024; Semenova 2024).
@@
The design-based model combines a model of randomization within an experiment with causal models of potential outcomes from Neyman (1923), Welch (1937), Kempthorne (1952), Copas (1973), Rubin (1974, 1977), Greenland and Robins (1986), Holland (1986) and others. Copas (1973) derives the design-based likelihood of a completely randomized experiment. We generalize the design-based model to derive the likelihood from a generic randomization process, which we demonstrate with the likelihood for Bernoulli randomization. While Copas (1973) uses the design-based likelihood to test hypotheses about the average effect and the Fisher null hypothesis, our advance is to use the likelihood to learn about a more primitive object: the joint distribution of potential outcomes in the sample, including the number of defiers. In doing so, we connect Copas’ design-based model of an experiment to the modern causal model, which reveals that experimental data contains information about the number of defiers in the sample, even if it does not contain information about their population analog. Likelihood-based analysis is uncommon in design-based models of experiments, partly due to its association with the sorts of restrictive “parametric” assumptions that experiments are constructed to avoid. One notable exception is Li and Ding (2016), who construct exact confidence intervals on the average effect using parametric distributions implied by the experiment design. By combining the potential outcomes framework with a parametric model of the randomization process itself, which is under the control of the designer of the experiment, the design-based model yields a likelihood function free of restrictive assumptions.
@@
Our use of a decision-theoretic rule in the style of Wald (1949), rather than a classical hypothesis test, facilitates our progress. Classical hypothesis tests specify a status-quo null hypothesis and place the burden of evidence on the alternative hypothesis. This asymmetric treatment of Type I and Type II errors can imply an unrealistically conservative decision maker (Tetenov, 2012) and exclude statistical evidence that, while weak, is still informative. To make use of all of the evidence present in our design-based likelihood, we focus on the statistical decision problem of choosing the correct joint distribution of potential outcomes in the sample, rather than on testing hypotheses about a prespecified distribution under prespecified size and power. We do not claim to provide a “test” of a monotonicity assumption. We provide a “decision” about the joint distribution of potential outcomes that can provide evidence for or against a monotonicity assumption.
@@
We contribute to the integration of statistical decision theory into econometrics (Manski, 2004; Dehejia, 2005; Manski, 2007; Hirano, 2008; Hirano and Porter, 2009; Stoye, 2012; Kitagawa and Tetenov, 2018; Manski, 2018, 2019; Manski and Tetenov, 2021; Fernández, Montiel Olea, Qiu, Stoye, and Tinda, 2024). Economics is a natural discipline to embrace statistical decision theory because statistical decision theory has such a tight relationship with economic theory. Statistical decision theory is natural to apply in our setting because it does not require large sample assumptions or asymptotic approximations. Our use of design-based statistical decision theory departs from “asymptotic analysis of statistical decision rules in econometrics” as reviewed in Hirano and Porter (2020) in favor of the literature that focuses on statistical decision theory in finite sample settings (Canner, 1970; Manski and Tetenov, 2007; Schlag, 2007; Stoye, 2007, 2009; Tetenov, 2012). Our auxiliary construction of credible sets on the joint distribution of potential outcomes in Table A.1 provides a Bayesian alternative to classical asymptotic inference on Fréchet bounds (Manski, Sandefur, McLanahan, and Powers, 1992; Horowitz and Manski, 2000; Tamer, Chernozhukov, and Hong, 2004) and the number of defiers within the Fréchet bounds (Imbens and Manski, 2004).
@@
Our proposed decision rule engages with justifications commonly employed in statistical decision theory. As we show in Figure 3, our decision varies systematically with the data. One challenge with statistical decision criteria such as minimax and minimax regret is that researchers have shown instances in which they yield rules that do not vary with the data (Schlag, 2003; Manski, 2004; Hirano and Porter, 2009; Stoye, 2009). In his influential introduction to statistical decision theory, Ferguson (1967) recognizes that because maximum likelihood rules vary with the data, they are “reasonable” because they are better than “just guessing.” “However,” he continues, “decision theory, as developed here, is devoted to the problem of finding optimal rules, so that we do not refer to maximum likelihood estimates […] unless they turn out naturally to be optimal in some sense.” He goes on to claim that “because of the basic theorem of decision theory,” which he refers to as the complete class theorem, “in general, every really good decision rule is practically Bayes.” In Section 2.4.2, we further justify our decision rule as Bayes optimal under the appropriate conditions. The Bayes optimality of our rule enables us to make important claims. Specifically, it implies that our rule is admissible—that is, there is no alternative rule that performs at least as well for all possible joint distributions of potential outcomes, and strictly better for some (Ferguson, 1967). Furthermore, it ensures that our rule cannot be bested in a betting framework (Freedman and Purves, 1969). It also allows us to quantify the relative losses from suboptimal rules, including a rule that places equal probability on all numbers of defiers within the Fréchet bounds, and a rule that assumes monotonicity, as we do in Section 2.4.2.
@@
In addition to standard decision theoretic justifications, we also offer a perspective that our design-based maximum likelihood decision rule is optimal through the principle of maximum entropy (Jaynes, 1957a, b), which does not require specification of a prior. Golan (2002) reviews a literature in econometrics that uses the principle of maximum entropy to abstract away from functional form assumptions on the likelihood function. In our setting, the design of the randomization, which is under the control of the experimenter, informs the functional form of the likelihood function, which also determines the functional form of the entropy.

### v4

@@ 2.4.3 The Maximum Likelihood Decision Rule Maximizes Entropy
We can also motivate our maximum likelihood rule through the principle of maximum entropy (jaynes1957a; jaynes1957b), which unites the theory of information with statistical physics. In Section 1.1, we discuss how differences in entropy drive variation in the likelihood such that the maximum likelihood distribution is also the maximum entropy distribution. jaynes1968 advocates for application of the principle of maximum entropy to circumvent the subjective specification of a prior. To make Bayesian decision-making more objective, one option is to choose the least informative prior. Jaynes’ alternative to make the process more objective is to choose the least informative updated distribution. The distribution that can generate the data in the greatest number of ways—the distribution that maximizes entropy, also the distribution that maximizes our design-based likelihood—is the least informative updated distribution. In the Monty Hall game, a celebrated application of Bayesian decision making, it is possible to make the same decision without specifying a Bayesian prior using a maximum likelihood decision rule and the principle of maximum entropy (wang2016).
3 Applications: Counting Defiers in Health Care
@@
We apply our rule to two published experiments with positive, statistically significant average effects on takeup of desired health behaviors and plausible defiers. Figure 3 summarizes the results from all possible empirical applications in experiments with given sizes, illustrating empirical conditions under which our rule yields defiers. It is predictable from the data for each application and the patterns in Figure 3 that our first application yields no defiers and our second application yields defiers, which we confirm with exhaustive grid search. Our focus here is to provide examples to illustrate how applied researchers can present and interpret quantitative results and to provide substantive insights. The top of Table 1 summarizes the context and design for both applications. The statistics in the table can be calculated using our Python- and Stata-compatible dbmle package, documented at https://pypi.org/project/dbmle/ (christy2025dbmle).
@@
In tappin2015, the sample includes 612 pregnant smokers, and the definition of “takeup” is to quit smoking, which is a desired health behavior given concerns that smoking can cause pregnancy complications and lasting adverse effects on children. The experiment uses a completely randomized design, with half of the subjects assigned to an intervention arm with a payment and the other half assigned to a control arm with usual care. Monotonicity assumptions have a long history in the context of interventions that induce pregnant women to quit smoking. In a study discussed by angrist1996, permutt1989 proposed an early monotonicity assumption requiring that women who would quit smoking in control would also quit smoking in an arm that received a multi-faceted intervention. However, literature in psychology and economics expresses concern that payments can backfire in many contexts, including on-time pickup of children from day care (gneezy2000), blood donation (mellstrom2008; lacetera2010), and vaccination (schneider2023). Defiers who would quit smoking in control but not in intervention are plausible in tappin2015 because a payment could backfire by crowding out the intrinsic motivations of pregnant women to quit smoking based on moral ideals and health goals.
@@
The average effect of the payment relative to usual care is a statistically significant 14 percentage point increase in the smoking quit rate on a base of 8% in control. Our proposed decision rule reveals evidence beyond the average effect: our maximum likelihood estimate includes $52/612=8\%$ always takers (effect=0), $86/612=14\%$ compliers (effect=1), and $474/612=77\%$ never takers (effect=0). The weighted average effect across the three types ($8\times 0+14\times 1+77\times 0$) is equal to the estimated average effect of 14 percentage points, which is also equal to the share of compliers. We would obtain the same results if we assumed monotonicity in the sample, so our rule provides evidence in favor of monotonicity without assuming it.
@@
We emphasize that even though our proposed decision rule provides evidence that there are no defiers in the sample, there is considerable uncertainty in our maximum likelihood estimate. As discussed in Appendix D, we report a 95% smallest credible set on the number of defiers in the sample, which we construct by reporting the collection of numbers of defiers occurring within the 95% smallest credible set on the joint distribution of potential outcomes in the sample. The smallest credible set on defiers includes up to 71 defiers, which represents 71/612=12% of the sample size.
Table 1: Standard Statistics and Proposed Design-Based Maximum Likelihood Estimates
@@
Building on the evidence provided by our decision rule, researchers could go on to make the stronger assumption of monotonicity and proceed as usual. With additional data on covariates, they could calculate average characteristics of always takers, compliers, and never takers (imbens1997; katz2001; abadie2002; abadie2003). With additional data on a second stage outcome like birth weight, they could interpret the resulting instrumental variable estimate as a local average treatment effect on compliers (imbens1994).
@@
Our second example, johnson2003, is a high-profile experiment in which a simple change from an opt-in default to an opt-out default increases organ donation almost twofold from 43% in control to 82% in intervention. Results from johnson2003 have inspired state policy changes to encourage life-saving organ donations to combat chronic organ shortages (kessler2025). However, in a recent experiment by kessler2025, in which takeup represents actually becoming an organ donor instead of hypothetically becoming an organ donor, the average effect is close to zero. Especially given the hypothetical outcome in johnson2003, both compliers and defiers are plausible. Compliers may think of the default as an endorsement by the experimenter and thus follow it regardless of what it is. Defiers may think of a default as a threat to their autonomy and thus go against it regardless of what it is.
@@
Given the plausibility of defiers in this context, a monotonicity assumption might not be reasonable, but there is no obvious alternative, so standard practice is to engage with bounds on defiers, which can be large. As shown at the bottom of Table E.1, in which we report auxiliary statistics, the smallest possible number of defiers in the sample is 0 because it is always possible for all people who take up to be always takers and all people who do not take up to be never takers, as in the null of Fisher’s exact test. The largest possible number of defiers in the sample occurs if all 11(=6150)11(=61-50) people who do not take up in intervention and all 23 people who take up in control are defiers, yielding a maximum defier share of $30\%(=(23+11)/115$). The estimated Fréchet bounds are tighter because they incorporate the estimated marginal distributions of potential outcomes in each arm using the observed takeup rates. To calculate the upper Fréchet bound from (3), we consider the bounds on the share of defiers implied by the takeup rates in intervention and control. The intervention takeup rate implies that the share of defiers in intervention and thus in the sample is at most $18\%(=100\%-82\%)$, and the control takeup rate implies that the share of defiers in control and thus in the sample is at most 43%, so the share in intervention binds, and the upper Fréchet bound is 18%.
@@
The Fréchet bounds are estimated with error, so there is even greater uncertainty in the number of defiers. The maximum number of defiers in our 95% smallest credible set is 34, which represents 30% of the sample of 115. The 95% confidence interval on the population fraction of defiers from imbens2004, incorporating the adjustment recommended by jun2023 with pretesting at the 0.01 level, also accounts for estimation of the Fréchet bounds, so we report it for comparison. It is slightly narrower, extending from 0% to 27%, but still includes the estimated Fréchet bounds. Accounting for estimation of the Fréchet bounds is empirically important. As demonstrated in Figure 2, if we do not account for estimation of the Fréchet bounds and instead construct a 95% smallest credible set on defiers within the estimated Fréchet set, variation in the likelihood allows us to exclude 9 defiers— a count in the middle of the Fréchet bounds translated into whole numbers from 0 to 21.
@@
In this application, evidence from the design-based model suggests a weaker alternative assumption to monotonicity—an assumption of no never takers—that can subsequently be imposed. Monotonicity can, of course, still be imposed with acknowledgment that it is a stronger assumption. While the assumption of no never takers is still strong, by the principle of maximum entropy, it “assumes the least” (jaynes1968) about the joint distribution of potential outcomes. From the design-based likelihood in (2), the assumption of no never takers can generate the data in $(28\text{ choose }15)\times(66\text{ choose }35)\times(21\text{ choose }11)=8.5\times 10^{31}$ randomized assignments representing 3.4% of the total randomized assignments ($(115\text{ choose }61)=2.5\times 10^{33}$). The assumption of no defiers can generate the data in only $(49\text{ choose }26)\times(45\text{ choose }24)\times(21\text{ choose }11)=7.8\times 10^{31}$ randomized assignments, representing 3.1% of the total randomized assignments. An ex post justification for the assumption of no never takers is that though experimental subjects might feel comfortable complying with or defying defaults imposed by the experimenter, they might feel uncomfortable revealing to the experimenter that they would not donate their organs under either default, especially since they do not know when they are asked a hypothetical question if a later question will ask them for a response under the other default. If an assumption of no never takers is not palatable, evidence for defiers could be used to motivate methods for examining effect heterogeneity, like the collection of additional data or the use of machine learning methods.
@@
If a second stage outcome such as actual organ donation is available, under an assumption of no never takers, the instrumental variable estimate is no longer interpretable as an average treatment effect on compliers without further assumptions. As discussed by angrist1996, under an additional assumption that average treatment effects are the same for compliers and defiers, the instrumental variable estimate is again interpretable as a LATE. dechaisemartin2017 and tchetgen2024 propose alternative assumptions to interpret the instrumental variable estimand in the presence of compliers and defiers.
@@
The value of an assumption of no never takers becomes apparent if data are available on covariates. Under the assumption of no never takers, it is possible to label some specific people as compliers and others as defiers: all people in intervention who do not take up must be compliers, and all people who do not take up in control must be defiers. This reasoning allows us to make claims that the intervention had a causal effect on some specific people in the sample. Using terminology from pearl1999, under an assumption of no never takers, the intervention was a “necessary” cause of no takeup for all 11 people assigned intervention who do not take up, and was a “sufficient” cause of no takeup for all 31 people assigned control who do not take up. Under a monotonicity assumption, it is not possible to directly label any specific people as compliers or defiers because people who take up in intervention can be compliers or always takers, and people who do not take up in control can be compliers and never takers. An assumption that allows us to label some specific people as compliers and others as defiers could be an important first step to learn their characteristics, improve targeting in future samples, and obtain a larger average effect.
@@ 4 Contributions to the Literature
Our primary insight is that a randomized experiment can reveal evidence on an estimand more primitive than the average effect: the joint distribution of potential outcomes in the sample. The evidence can support a monotonicity assumption or inform a specific alternative assumption. We obtain evidence using the structure of the implemented randomization process to construct a design-based causal model and its resulting likelihood function, and we evaluate this evidence in a statistical decision theory framework. Our work contributes to the literature on design-based econometrics, statistical decision theory, and causal inference without monotonicity.
@@
In his 1935 book on the design of experiments, Fisher notes that the design of Charles Darwin’s experiments was “greatly superior” to the “statistical methods available at the time” (fisher1935). In their 2017 handbook chapter on the econometrics of experiments, Susan Athey and Guido Imbens write that they “recommend using statistical methods that are directly justified by randomization, in contrast to the more traditional sampling-based approach that is commonly used in econometrics” (athey2017). The sampling-based approach assumes that people in an experiment are randomly sampled from a population, and it attempts to learn about the population. The design-based approach, which we adopt, attempts to learn about the sample of people in the experiment, which is useful since participation in experiments and clinical trials can be far from random (alsan2025) and some samples do not come from a meaningful population (abadie2020).
@@
Our use of a design-based model rather than a sampling-based model enables our contributions. Our ability to learn about defiers with a design-based likelihood may seem paradoxical because it is well known that the traditional sampling-based likelihood does not vary with the number of defiers within the Fréchet bounds determined by the estimated marginal distributions of potential outcomes in intervention and control. As we review in Appendix B, for a given pair of marginal distributions, the likelihood function in the traditional sampling-based model is flat over all copulas connecting them. A large literature has focused on specifying what we can learn from the nonparametric Fréchet bounds in the traditional sampling-based model (balke1997; heckman1997; manski1997mixing; tian2000; zhang2003; imbens2004; fan2010; mullahy2018; li2019; bai2024; semenova2024).
The design-based model allows us to construct estimates that vary within the Fréchet bounds, as in Figure 1, by harnessing the parametric structure that follows directly from the randomization process. Parametric structure within randomized experiments is easy to justify as long as it corresponds to the actual randomization process, which is under the control of the people who design and implement the experiment. In many applications, economists use the parameteric structure of a model to infer aggregate counterfactuals from observed individual-level heterogeneity. We invert this procedure: using aggregate counterfactuals observed from the experiment, we apply the parametric structure of the randomization process to infer individual-level heterogeneity.
@@
Our use of the likelihood implied by the design-based model is central to our contributions. The design-based model defines counterfactuals using causal models of potential outcomes from neyman1923, welch1937, kempthorne1952, copas1973, rubin1974; rubin1977, greenland1986, holland1986, and others, and then uses the randomization structure to derive the design-based likelihood. While copas1973 derives the design-based likelihood and uses it to test hypotheses about the average effect and the Fisher null hypothesis, our advance is to use the likelihood to learn about a more primitive object: the joint distribution of potential outcomes in the sample, including the number of defiers. We revisit Copas’ design-based model in light of modern causal models, modern assumptions about the absence of defiers, and modern computational advances. Likelihood-based models can provide exact distributions of experimental data, test statistics, and confidence intervals, but they remain uncommon relative to simulation methods in randomization inference. One exception is ding2019, who use the design-based likelihood to examine the sensitivity of inference on aggregate parameters to assumptions about the joint distribution of potential outcomes. rigdon2015 and li2016 construct exact confidence intervals without the design-based likelihood.
@@
Our use of a decision-theoretic rule in the style of wald1949, rather than a classical hypothesis test, facilitates our progress. Classical hypothesis tests specify a null hypothesis and place the burden of evidence on the alternative hypothesis. This asymmetric treatment of Type I and Type II errors can imply an unrealistically conservative decision maker (tetenov2012) and exclude statistical evidence that, while weak, is still informative. To make use of all the evidence in our design-based likelihood, we focus on the statistical decision problem of choosing the correct joint distribution of potential outcomes in the sample, rather than on testing hypotheses about a prespecified distribution under prespecified size and power. We do not claim to provide a “test” of a monotonicity assumption. We provide a “decision” about the joint distribution of potential outcomes that can provide evidence for or against a monotonicity assumption.
@@
Our use of design-based statistical decision theory contributes to the integration of statistical decision theory into econometrics (manski2004; dehejia2005; manski2007; hirano2008; hirano2009; stoye2012; kitagawa2018; manski2018; manski2019; manski2021; fernandez2024). Econometrics is a natural discipline to embrace statistical decision theory because statistical decision theory has such a tight relationship with economic theory. Statistical decision theory is natural to apply in our design-based setting because it does not require large sample assumptions or asymptotic approximations. Our use of design-based statistical decision theory departs from “asymptotic analysis of statistical decision rules in econometrics” as reviewed in hirano2020 in favor of the literature that focuses on statistical decision theory in finite sample settings (canner1970; manski2007admissible; schlag2007; stoye2007; stoye2009; tetenov2012).
@@
We justify our proposed decision rule with justifications commonly employed in statistical decision theory, starting with the justification that our decision varies systematically with the data, as we show in Figure 3. One challenge with statistical decision criteria such as minimax and minimax regret is that researchers have shown instances in which they yield rules that do not vary with the data (schlag2003; manski2004; hirano2009; stoye2009). In his influential introduction to statistical decision theory, ferguson1967 recognizes that because maximum likelihood rules vary with the data, they are “reasonable” because they are better than “just guessing.” “However,” he continues, “decision theory, as developed here, is devoted to the problem of finding optimal rules, so that we do not refer to maximum likelihood estimates […] unless they turn out naturally to be optimal in some sense.”
In Section 2.4.2, we further justify our decision rule as Bayes optimal under the appropriate conditions, which advances Bayesian decision theory (chamberlain2011) by incorporating the design-based likelihood for our class of experiments. Bayes optimality implies that our rule is admissible—that is, there is no alternative rule that performs at least as well for all possible joint distributions of potential outcomes, and strictly better for some (ferguson1967). Furthermore, it ensures that our rule cannot be bested in a betting framework (freedman1969). The Bayesian interpretation of our rule facilitates our auxiliary construction of credible sets on the joint distribution of potential outcomes in Table E.1, providing an exact Bayesian alternative to classical asymptotic inference on Fréchet bounds (manski1992; horowitz2000; tamer2004; jun2023) and the number of defiers within the Fréchet bounds (imbens2004).
@@
In addition to standard decision theoretic justifications, we also offer a perspective that our design-based maximum likelihood decision rule is optimal through the principle of maximum entropy (jaynes1957a; jaynes1957b), which does not require specification of a prior. golan2002 reviews a literature in econometrics that uses the principle of maximum entropy to abstract away from functional form assumptions on the likelihood function. In our setting, the design of the randomization, which is under the control of the experimenter, determines the functional form of the likelihood function, which also determines the functional form of the entropy.